# Text4Seg: Reimagining Image Segmentation as Text Generation

### v1

@@ 4.2 Referring Expression Segmentation
As summarized in Tab. 1, our Text4Seg achieves the highest performance across all splits of the refCOCO (+/g) datasets. For 7B-scale MLLMs, Text4Seg delivers an impressive average cIoU of 75.0, surpassing the closest competitor, Groundhog, which scores 74.2 cIoU. Notably, Text4Seg stands out with an average of 75.4 cIoU. At the 13B parameter scale, Text4Seg achieves a marked improvement, with an average cIoU of 76.2, significantly outperforming GSVA’s 72.8 cIoU. These results demonstrate the clear advantage of Text4Seg in single-object referring expression segmentation.
@@
| Methods | refCOCO | refCOCO+ | refCOCOg | Avg. | |||||
| val | testA | testB | val | testA | testB | val | test | ||
| Specialised Segmentation Models | |||||||||
| MDETR Kamath et al. (2021) | 86.8 | 89.6 | 81.4 | 79.5 | 84.1 | 70.6 | 81.6 | 80.9 | 81.8 |
| G-DINO Liu et al. (2023b) | 90.6 | 93.2 | 88.2 | 82.8 | 89.0 | 75.9 | 86.1 | 87.0 | 86.6 |
| UNINEXT-L Yan et al. (2023) | 91.4 | 93.7 | 88.9 | 83.1 | 87.9 | 76.2 | 86.9 | 87.5 | 87.0 |
| Generalist Segmentation Models (7B) | |||||||||
| Shikra Chen et al. (2023a) | 87.0 | 90.6 | 80.2 | 81.6 | 87.4 | 72.1 | 82.3 | 82.2 | 82.9 |
| Ferret You et al. (2023) | 87.5 | 91.4 | 82.5 | 80.8 | 87.4 | 73.1 | 83.9 | 84.8 | 83.9 |
| Qwen-VL Bai et al. (2023) | 88.6 | 92.3 | 84.5 | 82.8 | 88.6 | 76.8 | 86.0 | 86.3 | 85.7 |
| InternVL2-8B Chen et al. (2024) | 87.1 | 91.1 | 80.7 | 79.8 | 87.9 | 71.4 | 82.7 | 82.7 | 82.9 |
| LISA Lai et al. (2024) | 85.4 | 88.8 | 82.6 | 74.2 | 79.5 | 68.4 | 79.3 | 80.4 | 79.8 |
| GSVA Xia et al. (2024) | 86.3 | 89.2 | 83.8 | 72.8 | 78.8 | 68.0 | 81.6 | 81.8 | 80.3 |
| NEXT-Chat Zhang et al. (2023) | 85.5 | 90.0 | 77.9 | 77.2 | 84.5 | 68.0 | 80.1 | 79.8 | 80.4 |
| PixelLM Ren et al. (2024) | 89.8 | 92.2 | 86.4 | 83.2 | 87.0 | 78.9 | 84.6 | 86.0 | 86.0 |
| Groma Ma et al. (2024) | 89.5 | 92.1 | 86.3 | 83.9 | 88.9 | 78.1 | 86.4 | 87.0 | 86.5 |
| Text4Seg | 86.4 | 90.3 | 81.7 | 80.5 | 86.3 | 72.3 | 82.4 | 82.7 | 82.8 |
| Text4Seg | 89.6 | 93.3 | 85.4 | 84.2 | 90.2 | 78.5 | 84.4 | 84.7 | 86.3 |
| Text4Seg | 90.8 | 93.7 | 87.6 | 84.7 | 90.2 | 79.0 | 84.8 | 85.0 | 87.0 |
| Text4Seg | 89.7 | 93.0 | 85.8 | 84.6 | 90.1 | 78.6 | 85.0 | 85.1 | 86.5 |
| Text4Seg | 90.3 | 93.4 | 87.5 | 85.2 | 89.9 | 79.5 | 85.4 | 85.4 | 87.1 |
| Generalist Segmentation Models (13B) | |||||||||
| Shikra Chen et al. (2023a) | 87.8 | 91.1 | 81.8 | 82.9 | 87.8 | 74.4 | 82.6 | 83.2 | 84.0 |
| LISA Lai et al. (2024) | 85.9 | 89.1 | 83.2 | 74.9 | 81.1 | 68.9 | 80.1 | 81.5 | 80.6 |
| GSVA Xia et al. (2024) | 87.7 | 90.5 | 84.6 | 76.5 | 81.7 | 70.4 | 83.9 | 84.9 | 82.5 |
| Text4Seg | 91.2 | 94.3 | 88.0 | 85.7 | 90.8 | 80.1 | 85.6 | 85.5 | 87.7 |
@@ Results.
As shown in Tab. 3, our Text4Seg achieves the best results on the refCOCO and refCOCO+ datasets, while Groma performs well on refCOCOg. However, Text4Seg delivers the highest overall accuracy, reaching 87.1%. Notably, both Text4Seg and Text4Seg surpass their respective MLLM baselines. In particular, Text4Seg demonstrates a significant improvement over InternVL2-8B, increasing its average accuracy from 82.9% to 87.1%. Additionally, our Text4Seg outperforms previous SOTA, Shikra, by an average margin of 3.7%. These results highlight the superiority of our Text4Seg, which offers a finer, pixel-level representation that enhances the precision of bounding box predictions.
@@
Table 4 presents a comparison between LLaVA-1.5 and Text4Seg across various VQA and RES benchmarks. Text4Seg, trained on the mixed dataset, not only achieves performance comparable to LLaVA-1.5 in visual question answering tasks, but also demonstrates strong results in RES benchmarks. These results validate that our text generation based segmentation method acts as a seamless enhancement, offering a streamlined approach for pre-training MLLMs. It successfully integrates robust segmentation functionality without compromising the model’s conversational capabilities.
@@ 4.6 Ablation Study
Given that the focus of this work is on introducing semantic descriptors for visual segmentation and grounding, we conducted a series of ablation studies to assess the impact of semantic descriptors on performance, using InternVL2-8B Chen et al. (2024) as the MLLM.
@@
Tab. 6 compares the performance of various mask refiners, such as SAM with different architectures, against no refiner for semantic descriptors at a resolution. SAM with the ViT-L architecture achieves similar performance to SAM with ViT-H while reducing inference time. Notably, Text4Seg with ViT-L increases the average performance on RES tasks from 70.3 to 75.4 cIoU compared to Text4Seg without a mask refiner, with only a little increase in inference time.
@@ A.2 Details of Training Hyper-parameters
Appendix B Additional visual instruction data Details
@@
Appendix C Additional Quantitative Results
C.1 More Results on Mask Refiner
@@
C.2 More Results on Different Resolution of Semantic Descriptors.
@@
Appendix D Additional Qualitative Results

### v2

@@ 4.2 Referring Expression Segmentation
As summarized in Tab. 1, our Text4Seg achieves the highest performance across all splits of the refCOCO (+/g) datasets. For 7B-scale MLLMs, Text4Seg delivers an impressive average cIoU of 75.0, surpassing the closest competitor, Groundhog, which scores 74.2 cIoU. Notably, Text4Seg stands out with an average of 75.4 cIoU. At the 13B parameter scale, Text4Seg achieves a marked improvement, with an average cIoU of 76.2, significantly outperforming GSVA’s 72.8 cIoU. Even without using the SAM refiner, our method remains competitive. For instance, Text4Seg†, refined with CRFs, and Text4Seg‡, based on semantic descriptors, achieve results that rival or exceed existing methods.
@@
Methods | LLM | refCOCO | refCOCO+ | refCOCOg | Avg. | |||||
val | testA | testB | val | testA | testB | val | test | |||
Specialised Segmentation Models | ||||||||||
MDETR Kamath et al. (2021) | 86.8 | 89.6 | 81.4 | 79.5 | 84.1 | 70.6 | 81.6 | 80.9 | 81.8 | |
G-DINO Liu et al. (2023c) | 90.6 | 93.2 | 88.2 | 82.8 | 89.0 | 75.9 | 86.1 | 87.0 | 86.6 | |
PolyFormer-L Liu et al. (2023b) | 90.4 | 92.9 | 87.2 | 85.0 | 89.8 | 78.0 | 85.8 | 85.9 | 86.9 | |
UNINEXT-L Yan et al. (2023) | 91.4 | 93.7 | 88.9 | 83.1 | 87.9 | 76.2 | 86.9 | 87.5 | 87.0 | |
Generalist Segmentation Models (8B) | ||||||||||
Shikra Chen et al. (2023a) | Vicuna-7B | 87.0 | 90.6 | 80.2 | 81.6 | 87.4 | 72.1 | 82.3 | 82.2 | 82.9 |
Ferret You et al. (2023) | Vicuna-7B | 87.5 | 91.4 | 82.5 | 80.8 | 87.4 | 73.1 | 83.9 | 84.8 | 83.9 |
Qwen-VL Bai et al. (2023) | Qwen-7B | 88.6 | 92.3 | 84.5 | 82.8 | 88.6 | 76.8 | 86.0 | 86.3 | 85.7 |
InternVL2-8B Chen et al. (2024) | InternLM2.5-7B | 87.1 | 91.1 | 80.7 | 79.8 | 87.9 | 71.4 | 82.7 | 82.7 | 82.9 |
LISA Lai et al. (2024) | Vicuna-7B | 85.4 | 88.8 | 82.6 | 74.2 | 79.5 | 68.4 | 79.3 | 80.4 | 79.8 |
GSVA Xia et al. (2024) | Vicuna-7B | 86.3 | 89.2 | 83.8 | 72.8 | 78.8 | 68.0 | 81.6 | 81.8 | 80.3 |
NEXT-Chat Zhang et al. (2023) | Vicuna-7B | 85.5 | 90.0 | 77.9 | 77.2 | 84.5 | 68.0 | 80.1 | 79.8 | 80.4 |
PixelLM Ren et al. (2024) | Vicuna-7B | 89.8 | 92.2 | 86.4 | 83.2 | 87.0 | 78.9 | 84.6 | 86.0 | 86.0 |
Groma Ma et al. (2024) | Vicuna-7B | 89.5 | 92.1 | 86.3 | 83.9 | 88.9 | 78.1 | 86.4 | 87.0 | 86.5 |
Text4Seg | DeepSeek-1.3B | 86.4 | 90.3 | 81.7 | 80.5 | 86.3 | 72.3 | 82.4 | 82.7 | 82.8 |
Text4Seg∗ | DeepSeek-7B | 87.2 | 90.8 | 83.4 | 82.1 | 88.1 | 76.8 | 81.1 | 81.0 | 83.8 |
Text4Seg | DeepSeek-7B | 89.6 | 93.3 | 85.4 | 84.2 | 90.2 | 78.5 | 84.4 | 84.7 | 86.3 |
Text4Seg∗ | Qwen-7B | 87.2 | 90.1 | 83.6 | 82.1 | 87.4 | 76.6 | 81.5 | 81.3 | 83.7 |
Text4Seg | Qwen-7B | 89.7 | 93.0 | 85.8 | 84.6 | 90.1 | 78.6 | 85.0 | 85.1 | 86.5 |
Text4Seg∗ | Vicuna-7B | 89.2 | 92.0 | 86.4 | 83.4 | 88.6 | 78.0 | 81.7 | 82.4 | 85.2 |
Text4Seg | Vicuna-7B | 90.8 | 93.7 | 87.6 | 84.7 | 90.2 | 79.0 | 84.8 | 85.0 | 87.0 |
Text4Seg∗ | InternLM2.5-7B | 88.3 | 91.4 | 85.8 | 83.5 | 88.2 | 77.9 | 82.4 | 82.5 | 85.0 |
Text4Seg | InternLM2.5-7B | 90.3 | 93.4 | 87.5 | 85.2 | 89.9 | 79.5 | 85.4 | 85.4 | 87.1 |
Generalist Segmentation Models (13B) | ||||||||||
Shikra Chen et al. (2023a) | Vicuna-13B | 87.8 | 91.1 | 81.8 | 82.9 | 87.8 | 74.4 | 82.6 | 83.2 | 84.0 |
LISA Lai et al. (2024) | Vicuna-13B | 85.9 | 89.1 | 83.2 | 74.9 | 81.1 | 68.9 | 80.1 | 81.5 | 80.6 |
GSVA Xia et al. (2024) | Vicuna-13B | 87.7 | 90.5 | 84.6 | 76.5 | 81.7 | 70.4 | 83.9 | 84.9 | 82.5 |
Text4Seg∗ | Vicuna-13B | 89.6 | 92.3 | 87.0 | 84.4 | 89.0 | 79.1 | 82.9 | 82.9 | 85.9 |
Text4Seg | Vicuna-13B | 91.2 | 94.3 | 88.0 | 85.7 | 90.8 | 80.1 | 85.6 | 85.5 | 87.7 |
@@ Results.
As shown in Tab. 3, our Text4Seg achieves the best results on the refCOCO and refCOCO+ datasets, while Groma performs well on refCOCOg. However, Text4Seg delivers the highest overall accuracy, reaching 87.1%. Notably, both Text4Seg and Text4Seg surpass their respective MLLM baselines. In particular, Text4Seg demonstrates a significant improvement over InternVL2-8B, increasing its average accuracy from 82.9% to 87.1%. Additionally, our Text4Seg outperforms previous SOTA, Shikra, by an average margin of 3.7%. It is worth noting that Text4Seg∗ and Text4Seg∗, without a mask refiner, outperform their respective baseline counterparts. These results emphasize the superiority of Text4Seg in following instructions, leading to enhanced visual grounding ability.
@@
Table 4 presents a comparison between LLaVA-1.5 and Text4Seg across various VQA and RES benchmarks. Notably, Text4Seg, trained on a mixed dataset, achieves performance on par with LLaVA-1.5 in visual question answering tasks while delivering strong results in RES benchmarks. These results validate that our text generation based segmentation method acts as a seamless enhancement, offering a streamlined approach for pre-training MLLMs. It successfully integrates robust segmentation functionality without compromising the model’s conversational capabilities.
@@ 4.6 Ablation Study
Focusing on semantic descriptors for visual segmentation and grounding, we conducted ablation studies to evaluate its impact on performance using InternVL2-8B Chen et al. (2024) as the MLLM.
@@
Tab. 6 compares the performance of various mask refiners, such as SAM with different architectures, against no refiner for semantic descriptors at a resolution. SAM with the ViT-L architecture achieves similar performance to SAM with ViT-H while reducing inference time. Notably, Text4Seg with SAM-L increases the average performance on RES tasks from 73.5 to 79.1 cIoU compared to Text4Seg without a mask refiner, with only a little increase in inference time.
@@ 5 Conclusion
Acknowledgment.
This study is supported under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
@@ A.2 Details of Training Hyper-parameters
Appendix B Comparison of training datasets
Most prior methods follow a two-stage training paradigm: Continued Pre-Training (CPT) using large datasets, followed by Supervised Fine-Tuning (SFT) for specific tasks. The datasets used in these approaches are summarized in the following tables:
We can note that:
1.
For CPT, previous methods rely heavily on large and diverse datasets, whereas our approach, Text4Seg, eliminates this requirement, demonstrating superior efficiency and effectiveness.
2.
For SFT, we ensure a fair comparison by following previous works and train on:
The train split of refCOCO series for RES and REC tasks.
The train split of grefCOCO for the GRES task.
Methods | Datasets |
LISA | ADE20K, COCO-Stuff, PACO-LVIS, PartImageNet, PASCAL-Part, refCLEF, refCOCO, refCOCO+, refCOCOg, LLaVA-v1.5-mix665k |
PixelLM | ADE20K, COCO-Stuff, PACO-LVIS, refCLEF, refCOCO, refCOCO+, refCOCOg, LLAVA-150k, multi-target reasoning segmentation (MUSE) |
GSVA | ADE20K, COCO-Stuff, PACO-LVIS, Mapillary Vistas, PASCAL-Part, refCLEF, refCOCO, refCOCO+, refCOCOg, gRefCOCO, LLaVA-Instruct-150K, ReasonSeg |
AnyRef | ADE20K, COCO-Stuff, PACO-LVIS, refCLEF, refCOCO, refCOCO+, refCOCOg, PhraseCut, Flickr30K Entities, AVSBench |
NEXT-Chat | Flickr30K Entities, Visual Genome, RefCOCO, RefCOCO+, RefCOCOg, VQAv2, PointQA, Visual7W, VCR, LLaVA-Instruct-150K, VG grounded captioning, Shikra-RD |
Groundhog | Multi-Modal Multi-Grained Grounding dataset (M3G2): PNG, Flickr30K-Entity, refCLEF, refCOCO, refCOCO+, refCOCOg, gRefCOCO, PhraseCut, D-Cube, ReasonSeg, RIO, SK-VG, VizWiz-G, TextVQA-X, GQA, VQS, Shikra-BinaryQA, EntityCount, FoodSeg-QA, LVIS-QA, RefCOCO-REG, RefCOCO+-REG, RefCOCOg-REG, gRefCOCO-REG, VG-SpotCap, V7W, PointQA, VCR, ShikraRD, SVIT-RD, Guesswhat, VG-RefMatch, HierText |
GLaMM | Grounding-anything Dataset (GranD): 11M images, 810M masks, 84M referring expressions, GranD-f |
Text4Seg | None |
Methods | Datasets |
LISA | refCOCO, refCOCO+, refCOCOg |
PixelLM | None |
GSVA | refCOCO, refCOCO+, refCOCOg |
AnyRef | refCOCO, refCOCO+, refCOCOg |
NEXT-Chat | refCOCO, refCOCO+, refCOCOg |
Groundhog | None |
GLaMM | refCOCO, refCOCO+, refCOCOg |
Text4Seg | refCOCO, refCOCO+, refCOCOg, refCLEF† |
Methods | Datasets |
LISA | grefCOCO |
GSVA | grefCOCO |
Text4Seg | grefCOCO |
Appendix C Additional visual instruction data Details
@@
Appendix D Additional Quantitative Results
D.1 More Results on Mask Refiner
@@
D.2 More Results on Different Resolution of Semantic Descriptors
@@
D.3 More Results regarding the Mask Refiner
We provide additional quantitative results on Tabs. 12, 13 and 14. While Text4Seg without a mask refiner slightly lags behind LISA and GSVA in terms of average cIoU on referring expression segmentation (RES) tasks, traditional mask refinement techniques, such as Conditional Random Fields (CRF), can be employed to enhance segmentation accuracy. For instance, Text4Seg with a CRF refiner improves the baseline performance from an average cIoU of 67.5 to 70.1 on RES tasks. Additionally, when using semantic descriptors, Text4Seg outperforms its counterpart with descriptors. Specifically, Text4Seg with semantic descriptors achieves an average cIoU of 71.4, surpassing LISA’s 69.9 and matching GSVA’s 71.4 on RES tasks. On the GRES tasks, as shown in the Tab. 13, both CRF and SAM refiners significantly enhance performance, outperforming LISA and GSVA. Notably, Text4Seg with semantic descriptors, even without a mask refiner, achieves performance superior to existing methods. Finally, on the REC tasks, Text4Seg without a SAM refiner continues to outperform current methods, further demonstrating the effectiveness of Text4Seg’s visual grounding capabilities.
Methods | Refiner | refCOCO | refCOCO+ | refCOCOg | Avg. | |||||
val | testA | testB | val | testA | testB | val | test | |||
Generalist Segmentation Models (8B) | ||||||||||
LISA Lai et al. (2024) | - | 74.9 | 79.1 | 72.3 | 65.1 | 70.8 | 58.1 | 67.9 | 70.6 | 69.9 |
GSVA Xia et al. (2024) | - | 77.2 | 78.9 | 73.5 | 65.9 | 69.6 | 59.8 | 72.7 | 73.3 | 71.4 |
Text4Seg | None | 66.2 | 68.7 | 63.6 | 60.7 | 64.5 | 54.9 | 64.2 | 64.2 | 63.4 |
Text4Seg | SAM-H | 75.0 | 78.6 | 70.1 | 68.4 | 73.4 | 60.0 | 71.5 | 71.7 | 71.1 |
Text4Seg | None | 69.7 | 71.2 | 67.9 | 64.5 | 68.0 | 60.2 | 66.6 | 66.7 | 66.9 |
Text4Seg | CRF | 72.6 | 74.8 | 70.0 | 67.2 | 71.5 | 62.2 | 69.1 | 69.4 | 69.6 |
Text4Seg | SAM-H | 78.8 | 81.5 | 74.9 | 72.5 | 77.4 | 65.9 | 74.3 | 74.4 | 75.0 |
Text4Seg | None | 68.3 | 70.0 | 67.3 | 63.1 | 67.2 | 59.9 | 66.5 | 66.4 | 66.1 |
Text4Seg | CRF | 71.3 | 73.7 | 69.6 | 65.9 | 70.4 | 61.9 | 69.3 | 69.3 | 68.9 |
Text4Seg | SAM-H | 78.0 | 80.9 | 74.6 | 71.6 | 77.3 | 66.0 | 74.8 | 74.7 | 74.7 |
Text4Seg | None | 70.5 | 72.3 | 69.3 | 64.4 | 68.7 | 60.6 | 65.1 | 66.5 | 67.2 |
Text4Seg | CRF | 73.2 | 75.7 | 71.4 | 67.0 | 71.9 | 62.4 | 67.3 | 68.9 | 69.7 |
Text4Seg | SAM-H | 79.3 | 81.9 | 76.2 | 72.1 | 77.6 | 66.1 | 72.1 | 73.9 | 74.9 |
Text4Seg | None | 70.3 | 71.9 | 68.7 | 65.0 | 68.9 | 60.8 | 66.7 | 67.6 | 67.5 |
Text4Seg | CRF | 73.0 | 75.2 | 70.7 | 67.6 | 72.1 | 62.6 | 68.9 | 70.3 | 70.1 |
Text4Seg | SAM-H | 79.2 | 81.7 | 75.6 | 72.8 | 77.9 | 66.5 | 74.0 | 75.3 | 75.4 |
Text4Seg‡ | None | 74.7 | 77.4 | 71.6 | 68.5 | 73.6 | 62.9 | 70.7 | 71.6 | 71.4 |
Text4Seg‡ | SAM-H | 78.6 | 81.7 | 74.3 | 71.8 | 77.4 | 65.1 | 73.9 | 74.7 | 74.7 |
Generalist Segmentation Models (13B) | ||||||||||
LISA Lai et al. (2024) | - | 76.0 | 78.8 | 72.9 | 65.0 | 70.2 | 58.1 | 69.5 | 70.5 | 70.1 |
GSVA Xia et al. (2024) | - | 78.2 | 80.4 | 74.2 | 67.4 | 71.5 | 60.9 | 74.2 | 75.6 | 72.8 |
Text4Seg | None | 71.3 | 72.9 | 70.3 | 65.9 | 70.0 | 61.8 | 66.8 | 67.6 | 68.3 |
Text4Seg | CRF | 74.1 | 76.4 | 72.4 | 68.5 | 72.8 | 63.6 | 69.1 | 70.1 | 70.9 |
Text4Seg | SAM-H | 80.2 | 82.7 | 77.3 | 73.7 | 78.6 | 67.6 | 74.0 | 75.1 | 76.2 |
Methods | Refiner | Validation Set | Test Set A | Test Set B | Avg. | |||
gIoU | cIoU | gIoU | cIoU | gIoU | cIoU | |||
Generalist Segmentation Models (8B) | ||||||||
LISA Lai et al. (2024) | - | 61.6 | 61.8 | 66.3 | 68.5 | 58.8 | 60.6 | 62.9 |
GSVA Xia et al. (2024) | - | 66.5 | 63.3 | 71.1 | 69.9 | 62.2 | 60.5 | 65.6 |
Text4Seg | None | 64.3 | 57.2 | 62.2 | 61.2 | 57.1 | 54.9 | 59.5 |
Text4Seg | SAM-H | 69.9 | 63.2 | 69.7 | 67.5 | 62.3 | 59.8 | 65.4 |
Text4Seg | None | 69.0 | 62.7 | 66.3 | 65.9 | 62.1 | 61.1 | 64.5 |
Text4Seg | CRF | 70.4 | 65.8 | 68.9 | 69.9 | 63.2 | 63.6 | 67.0 |
Text4Seg | SAM-H | 74.7 | 69.0 | 74.3 | 73.0 | 67.4 | 66.3 | 70.8 |
Text4Seg | None | 68.5 | 61.1 | 64.6 | 63.6 | 61.1 | 59.6 | 63.1 |
Text4Seg | CRF | 69.7 | 64.1 | 67.4 | 67.8 | 62.4 | 62.3 | 65.6 |
Text4Seg | SAM-H | 74.4 | 68.1 | 73.1 | 71.5 | 66.7 | 65.3 | 69.9 |
Text4Seg | None | 67.9 | 61.6 | 66.2 | 65.9 | 60.9 | 59.8 | 63.7 |
Text4Seg | CRF | 69.1 | 64.7 | 69.9 | 70.8 | 62.1 | 62.3 | 66.5 |
Text4Seg | SAM-H | 73.6 | 67.9 | 74.1 | 72.8 | 66.1 | 64.8 | 69.9 |
Text4Seg | None | 68.8 | 63.1 | 66.9 | 67.1 | 62.1 | 61.6 | 64.9 |
Text4Seg | CRF | 70.0 | 66.1 | 69.4 | 70.9 | 63.1 | 64.1 | 67.3 |
Text4Seg | SAM-H | 74.4 | 69.1 | 75.1 | 73.8 | 67.3 | 66.6 | 71.1 |
Text4Seg‡ | None | 71.8 | 65.6 | 71.2 | 70.0 | 64.2 | 62.5 | 67.6 |
Text4Seg‡ | SAM-H | 74.9 | 68.8 | 75.4 | 73.6 | 67.0 | 65.1 | 70.8 |
Generalist Segmentation Models (13B) | ||||||||
LISA Lai et al. (2024) | - | 63.5 | 63.0 | 68.2 | 69.7 | 61.8 | 62.2 | 64.7 |
GSVA Xia et al. (2024) | - | 68.0 | 64.1 | 71.8 | 70.5 | 63.8 | 61.3 | 66.6 |
Text4Seg | None | 69.2 | 63.9 | 67.4 | 67.6 | 62.7 | 62.0 | 65.5 |
Text4Seg | CRF | 70.3 | 66.9 | 69.8 | 71.4 | 63.8 | 64.4 | 67.8 |
Text4Seg | SAM-H | 74.8 | 69.8 | 75.1 | 74.3 | 68.0 | 67.1 | 71.5 |
Methods | Refiner | refCOCO | refCOCO+ | refCOCOg | Avg. | |||||
val | testA | testB | val | testA | testB | val | test | |||
Generalist Segmentation Models (8B) | ||||||||||
LISA Lai et al. (2024) | - | 85.4 | 88.8 | 82.6 | 74.2 | 79.5 | 68.4 | 79.3 | 80.4 | 79.8 |
GSVA Xia et al. (2024) | - | 86.3 | 89.2 | 83.8 | 72.8 | 78.8 | 68.0 | 81.6 | 81.8 | 80.3 |
Text4Seg | None | 83.6 | 87.3 | 79.1 | 78.0 | 83.6 | 70.3 | 78.5 | 78.8 | 79.9 |
Text4Seg | SAM-H | 86.4 | 90.3 | 81.7 | 80.5 | 86.3 | 72.3 | 82.4 | 82.7 | 82.8 |
Text4Seg | None | 87.2 | 90.8 | 83.4 | 82.1 | 88.1 | 76.8 | 81.1 | 81.0 | 83.8 |
Text4Seg | SAM-H | 89.6 | 93.3 | 85.4 | 84.2 | 90.2 | 78.5 | 84.4 | 84.7 | 86.3 |
Text4Seg | None | 87.2 | 90.1 | 83.6 | 82.1 | 87.4 | 76.6 | 81.5 | 81.3 | 83.7 |
Text4Seg | SAM-H | 89.7 | 93.0 | 85.8 | 84.6 | 90.1 | 78.6 | 85.0 | 85.1 | 86.5 |
Text4Seg | None | 89.2 | 92.0 | 86.4 | 83.4 | 88.6 | 78.0 | 81.7 | 82.4 | 85.2 |
Text4Seg | SAM-H | 90.8 | 93.7 | 87.6 | 84.7 | 90.2 | 79.0 | 84.8 | 85.0 | 87.0 |
Text4Seg | None | 88.3 | 91.4 | 85.8 | 83.5 | 88.2 | 77.9 | 82.4 | 82.5 | 85.0 |
Text4Seg | SAM-H | 90.3 | 93.4 | 87.5 | 85.2 | 89.9 | 79.5 | 85.4 | 85.4 | 87.1 |
Text4Seg‡ | None | 88.9 | 92.4 | 84.1 | 83.1 | 88.6 | 77.3 | 83.6 | 83.8 | 85.2 |
Text4Seg‡ | SAM-H | 89.6 | 92.6 | 84.9 | 83.7 | 88.8 | 77.6 | 84.6 | 84.8 | 85.8 |
Generalist Segmentation Models (13B) | ||||||||||
Shikra Chen et al. (2023a) | Vicuna-13B | 87.8 | 91.1 | 81.8 | 82.9 | 87.8 | 74.4 | 82.6 | 83.2 | 84.0 |
LISA Lai et al. (2024) | - | 85.9 | 89.1 | 83.2 | 74.9 | 81.1 | 68.9 | 80.1 | 81.5 | 80.6 |
GSVA Xia et al. (2024) | - | 87.7 | 90.5 | 84.6 | 76.5 | 81.7 | 70.4 | 83.9 | 84.9 | 82.5 |
Text4Seg | None | 89.6 | 92.3 | 87.0 | 84.4 | 89.0 | 79.1 | 82.9 | 82.9 | 85.9 |
Text4Seg | SAM-H | 91.2 | 94.3 | 88.0 | 85.7 | 90.8 | 80.1 | 85.6 | 85.5 | 87.7 |
Appendix E Additional Qualitative Results